What is the width of a rectangle that has an area of 1/6 cm2 and a legnth of 1.5 cm?
Found 2 solutions by Timnewman, Alan3354:
Answer by Timnewman(323)

hi!
Recall that:
Area (A) of a rectangle is given by
A=Lenght*Width
(that is) A=LW
but A=1/6cm²,L=1.5cm
therefore,
1.5w=1/6
1.5w=0.17
w=0.17/1.5
w=0.1cm
The width of the rectangle is 0.1cm

Answer by Alan3354(69443)

What is the width of a rectangle that has an area of 1/6 cm2 and a legnth of 1.5 cm?
==============
1/6 = 1.5*W
1/3 = 3W
W = 1/9 cm